OpenStudy (anonymous):

In this puzzle, each letter represents a different digit. What does the O represent and what does the E Represent? O +O=E O x O=E M x M=OM

OpenStudy (kinggeorge):

Since \(O+O=O \times O\), we know \(2O=O^2\) so \(O=2\) and \(E=4\)

OpenStudy (rulnick):

so M=5

OpenStudy (anonymous):

thankyou . Those Always Get me stuck I suck at math big time and the tutoring room is closed right now.

OpenStudy (kinggeorge):

you're welcome.